OpenStudy (anonymous):

Jennifer can ride her bike at 20 miles per hour. What is her speed in miles per second?

OpenStudy (anonymous):

Hi, I would start by thinking if one hour is equal to 20 miles, then 60 minutes equals 20 miles, so 1 minute would be 20/60 miles. Then repeat to find miles per second.

OpenStudy (carson889):

Its a game of fractions. First we state that we know in one hour there are 60 minutes. Thus as a statement we can write this as: 1 hour / 60 minutes , which directly translates to 1 hour PER 60 minutes. The division symbol also means PER. That is why miles per hour is written formally as m/hr and abbreviated MPH. So after multiplying our fraction of 20 miles / hour by our known value of 1 hour / 60 minutes (cross out the hour on the bottom of 20 miles and cross out hour on the top of 60 minutes) we get 1 mile / 3 minutes. Take this number and multiply by the fraction of 1 minute / 60 seconds. Doing this we find 1/180 miles per second. 0.0056 miles per second. \[\frac{ 20miles }{ hour }*\frac{ 1hour }{ 60minutes } = \frac{ 1mile }{ 3 minutes }\] \[\frac{ 1mile }{ 3minutes }*\frac{ 1 minute }{ 60seconds } = \frac{ 1 }{ 180 }\frac{ miles }{ second }\]

OpenStudy (anonymous):

A fact worth memorizing is that 60mph = 88 feet/sec. Solve the following for x: \[\frac{20}{60}=\frac{x}{88}\]\[x=\frac{88}{3} \text{feet}/\text{second} \]\[\frac{\frac{88}{3}}{5280}=\frac{1}{180}=0.00555556 \text{ miles}/\text{second} \]